Please help? I am having trouble with mixture formulas? How many kg of pure silver must be added to 56.75 kg of an alloy containing 9 percent to raise it to 15 percent?
Answer by checkley79(3341)

.09*56.75+X=.15(56.75+X)
5.1075+X=8.5125+.15X
X-.15X=8.5125-5.1075
.85X=3.405
X=3.405/.85
X=4.006 kg. OF PURE SILVER IS USED.
PROOF:
.09*56.75+4.006=.15(56.75+4.006
5.1075+4.006=.15(56.75+4.006)
9.1135=.15*60.756
9.1135=9.1134